OpenStudy (anonymous):

rationalize then simplify the answer: 4(square root of 5) / 3

OpenStudy (anonymous):

Multiply the top and bottom by \[4\sqrt{5}\]

OpenStudy (anonymous):

\[\frac{4\sqrt{5}}{3}\] is already rationalized

OpenStudy (anonymous):

didnt say if you wanted the numerator or denomiator rationalized

OpenStudy (anonymous):

numerator

OpenStudy (anonymous):

exactly then multiply top and bottome by 4sqrt5

OpenStudy (anonymous):

then I would get 80/12sqrt5

OpenStudy (anonymous):

should get :\[(80)/12\sqrt{5}\]

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

oooh got it. you can get away with just multiplying by \[\frac{\sqrt{5}}{\sqrt{5}}\] though right?